# Destruction of ’Peas in a Pod?’ A Candidate Multi-planet System Around the Nearby, Bright Star, HD208487

### v1

@@ Abstract
Context. The search for extrasolar planets (aka exoplanets) is continuously intensifying as new and improved instrumentation provides more precise data. Discovering new exoplanets orbiting bright stars in the local solar neighborhood is particularly important, as they tend to provide the best targets for future follow-up studies.
Aims. We re-investigate the HD208487 star system to test the reality of the proposed HD208487c world. We also search for additional companions using applied Bayesian statistics and 15+ years of new radial velocity (RV) data from the HARPS and the PFS instruments that were taken post-discovery of HD208487b, as part of our continued study of bright Sun-like stars within 50 pc of the Sun.
Methods. The RV data was analyzed with Generalized Lomb-Scargle Periodograms, followed by Bayesian analysis techniques using the EMPEROR code. We also scrutinised various stellar activity indices to search for any corresponding peaks in the power spectra, correlations with the RV measurements, or significant signals from a Bayesian analysis methodology. Finally, photometric data was also checked to test for any transits or possible activity manifestations that could lead to possible false RV signals or excess noise.
@@
Conclusions. We have uncovered a candidate three planet system that would consist of an inner gas giant, a central Saturn and an outer super-Neptune/sub-Saturn. Extensive analysis of both photometric and spectroscopic data as activity proxies strongly supports the planetary system hypothesis, yet more long-term RV data would help add more statistical weight to the reality of candidate planets c and d.Assuming our model best represents reality, a dynamical analysis suggests that gravitational scattering of an initially ordered, equally-spaced system in a long resonant chain of six Neptunes can explain the current architecture of HD208487- a moderately-eccentric, inner massive planet with an outer nearly resonant ($P_{d}/P_{c}=1.495$) gas giant and super-Neptune. More RVs may also shed light on the reality of a fourth Doppler signal uncovered in the data that sits close to the 2:1 period-ratio with signal of HD208487c.
@@
Exoplanets – HD208487 – Neptune – Multiplanetary Systems – Radial Velocity Method
@@ 1.1 Planet System Configurations
In this paper, we will discuss how our new analysis has lead to the statistical confirmation and discovery of two new planets in this system, HD208487c and HD208487d. The rest of the paper will follow the ensuing structure: §2 will discuss the RV observations in detail, including a breakdown of the measurements by telescope. In §3 we present updated stellar parameters for HD208487 using the SPECIES and ARIADNE codes. We begin the analysis of the RV dataset by observing the GLS periodograms in §4, while §5 details the detection of the three planets with the more rigorous Bayesian analysis. Next, we highlight the search for correlations between stellar activity indices and the RVs in §6. §7 describes the analysis of ASAS photometric data and the fitting of stellar cycles. Finally, we discuss the results and their implications in §8, followed by the conclusions in §9.
@@ 2.2.2 TERRA2
2.3 6.5m Carnegie Planet Finder Spectograph (PFS)
@@ 4.2 Split RV Data Set
As a first check to test how stellar activity is adding noise to the RV data, or indeed causing the signals we are seeing, we repeated the periodogram analysis with split data sets, where the split is made after calculating the median S-index value in each data set (excluding UCLES data, which do not have S-index values recorded). We then split the RVs in half: one data set with all RVs corresponding to S-index values less than the median and one data set with all RVs corresponding to S-index values greater than or equal to the median. We found that the low S-index data finds all three signals, while the high S-index data only finds the inner gas giant. Thus, the low S-index data provides a significantly better fit to the data when compared to the high S-index data, which would be expected if the RV signals are all independent of the stellar activity.
@@ 5 Bayesian Analysis
After examining the results from the periodogram analyses, we found significant evidence for two additional planetary signals in the data, along with the previously confirmed planet, HD208487b. In order to corroborate the initial results, we followed similar procedures to those employed in Tuomi et al. (2014); Jenkins & Tuomi (2014); D´ıaz et al. (2020) to model the RVs. The Markov Chain Monte Carlo (MCMC) code, Exoplanet Mcmc Parallel tEmpering Radial velOcity fitteR [^9] https://github.com/ReddTea/astroemperor/tree/master (EMPEROR  Peña & Jenkins, in preparation), ”searches for signals in radial velocity timeseries, employing Adaptive Parallel Tempering MCMC, convergence tests and Bayesian statistics, along with various noise models.” Further, the Python code provides Bayesian posterior distribution models and the parameterization for the planetary signals to describe the best-fit orbital solution to the star system. An early version of EMPEROR has already been used in a number of works to detect small signals in RV data (Wittenmyer et al., 2017; Jenkins et al., 2019a, b; Barnes et al., 2020; Jenkins et al., 2020), where the updated version employs a custom-written Adaptive Parallel Tempering MCMC engine, similar to the EMCEE package (Foreman-Mackey et al., 2013), as used in (Gregory, 2005, 2007).
@@ 5.2 Model Selection
Additionally, we ran each of these models with a fixed eccentricity of zero, a tightly constrained eccentricity described by a normal distribution centered at zero with a standard deviation of 0.1, and a constrained eccentricity described by a normal distribution centered at zero with a standard deviation of 0.3 (as explained in Jenkins et al. 2025, submitted), which enables EMPEROR to explore a wide variety of solutions based on the potential planetary eccentricities. In general, we found that fixing the eccentricity prior to zero would cause additional signals to appear that were harmonics of the real signals, primarily a consequence of the fact that the largest signal in the data corresponds to a massive planet on a non-circular orbit, and hence is poorly explained by a strict cosine-like function. All circular models that caused the 1st harmonic of this primary signal (HD208487b) to appear as a secondary signal in the data performed statistically much worse, enabling us to securely disregard these harmonic models. On the other hand, most of the models with the constrained eccentricity prior or the tightly constrained eccentricity prior yield very similar results. The main difference between the constrained and tightly constrained models are the predicted eccentricities of the outer two planets. The periods and semi-amplitudes are similar, but the constrained models tend to terminate with higher eccentricities, which results in both worse final statistics and long-term unstable planetary dynamics. In the theoretical limit of infinite walkers and steps, we would expect all of the models to converge to the same parameter configuration, even though we extended our analyses as far as reasonably possible to 288 million iterations.
@@ 5.2.2 Comparison of Earlier Data and Later Data
$P$ (days) $129.33^{+0.04}_{-0.04}$ $129.32^{+0.03}_{-0.01}$ $129.39^{+0.15}_{-0.22}$ $129.34^{+0.01}_{-0.03}$ $129.27^{+0.23}_{-0.14}$ $\mathbf{129.36^{+0.02}_{-0.02}}$
$K$ (m/s) $18.38^{+0.08}_{-0.23}$ $18.76^{+0.27}_{-0.35}$ $17.35^{+0.95}_{-1.22}$ $18.53^{+0.10}_{-0.17}$ $16.70^{+0.95}_{-0.02}$ $\mathbf{18.28^{+0.29}_{-0.40}}$
$e$ $0.36^{+0.01}_{-0.02}$ $0.35^{+0.01}_{-0.01}$ $0.18^{+0.04}_{-0.06}$ $0.35^{+0.00}_{-0.01}$ $0.13^{+0.03}_{-0.02}$ $\mathbf{0.37^{+0.01}_{-0.01}}$
$\omega$ (rad) $1.21^{+0.03}_{-0.01}$ $1.21^{+0.01}_{-0.03}$ $41^{+0.55}_{-0.35} 1.1$$16^{+0.06}_{-0.04}$ $2.58^{+0.48}_{-0.65}$ $\mathbf{1.10^{+0.01}_{-0.01}}$
$M_{0}$ (rad) $3.43^{+0.08}_{-0.10}$ $3.40^{+0.07}_{-0.01}$ $1.83^{+0.38}_{-0.68}$ $3.23^{+0.04}_{-0.05}$ $0.58^{+0.78}_{-0.49}$ $\mathbf{1.95^{+0.05}_{-0.04}}$
$m\sin i$ ($M_{J}$) $0.47^{+0.01}_{-0.01}$ $0.48^{+0.01}_{-0.01}$ $0.47^{+0.03}_{-0.03}$ $0.47^{+0.01}_{-0.01}$ $0.45^{+0.03}_{-0.00}$ $\mathbf{0.46^{+0.01}_{-0.00}}$
$AU1$ $0.52^{+0.00}_{-0.00}$ $0.52^{+0.00}_{-0.00}$ $0.53^{+0.00}_{-0.00}$ $0.53^{+0.00}_{-0.00}$ $0.52^{+0.00}_{-0.00}$ $\mathbf{0.53^{+0.00}_{-0.00}}$
$P_{2}$ (days) $490.82^{+5.62}_{-0.31}$ $488.10^{+0.89}_{-1.54}$ $934.56^{+17.97}_{-42.29}$ $485.60^{+4.71}_{-6.68}$ $923.92^{+26.38}_{-19.30}$ $\mathbf{923.06^{+2.02}_{-2.76}}$
$K_{2}$ (m/s) $3.73^{+0.27}_{-0.01}$ $4.35^{+0.33}_{-0.46}$ $9.76^{+1.26}_{-1.98}$ $4.03^{+0.17}_{-0.09}$ $8.69^{+1.13}_{-0.21}$ $\mathbf{6.18^{+0.09}_{-0.07}}$
$e_{2}$ $0.17^{+0.02}_{-0.06}$ $0.30^{+0.06}_{-0.10}$ $0.17^{+0.06}_{-0.12}$ $22^{+0.01}_{-0.04} 0.0$$08^{+0.03}_{-0.03}$ $\mathbf{0.19^{+0.01}_{-0.02}}$
$\omega_{2}$ (rad) $4.55^{+0.01}_{-0.40}$ $4.72^{+0.33}_{-0.48}$ $3.89^{+1.20}_{-0.04}$ $4.00^{+0.14}_{-0.17}$ $2.94^{+1.46}_{-1.10}$ $\mathbf{5.96^{+0.30}_{-5.96}}$
$M_{02}$ (rad) $0.74^{+0.90}_{-0.98}$ $0.41^{+0.19}_{-0.14}$ $3.61^{+1.34}_{-1.70}$ $2.39^{+0.00}_{-0.32}$ $3.96^{+0.00}_{-1.07}$ $\mathbf{1.29^{+0.52}_{-0.29}}$
$m_{2}\sin i$ ($M_{J}$) $0.16^{+0.01}_{-0.01}$ $0.18^{+0.01}_{-0.01}$ $0.51^{+0.07}_{-0.07}$ $0.17^{+0.01}_{-0.01}$ $0.46^{+0.06}_{-0.01}$ $\mathbf{0.32^{+0.01}_{-0.01}}$
$AU2$ $1.28^{+0.01}_{-0.01}$ $1.27^{+0.01}_{-0.01}$ $1.96^{+0.03}_{-0.03}$ $1.27^{+0.01}_{-0.01}$ $1.95^{+0.04}_{-0.03}$ $\mathbf{1.94^{+0.01}_{-0.01}}$
$P_{3}$ (days) $959.17^{+419.38}_{-194.63}$ $941.38^{+28.27}_{-18.51}$ $1294.65^{+1191.79}_{-38.62}$ $979.74^{+33.79}_{-36.32}$ $186.57^{+1125.53}_{-186.57}$ $\mathbf{1380.13^{+19.20}_{-8.25}}$
$K_{3}$ (m/s) $3.16^{+1.01}_{-0.94}$ $3.54^{+0.36}_{-0.50}$ $5.61^{+1.63}_{-3.77}$ $3.09^{+0.42}_{-0.37}$ $4.35^{+0.07}_{-1.57}$ $\mathbf{2.46^{+0.30}_{-0.09}}$
[…]
$\omega_{3}$ (rad) $1.08^{+2.25}_{-1.69}$ $0.57^{+2.24}_{-1.65}$ $2.80^{+1.49}_{-0.65}$ $3.34^{+0.12}_{-1.12}$ $4.91^{+1.86}_{-3.09}$ $\mathbf{0.29^{+0.28}_{-0.09}}$
$M_{03}$ (rad) $3.30^{+0.80}_{-1.37}$ $3.27^{+0.37}_{-0.82}$ $5.47^{+1.85}_{-3.32}$ $5.67^{+2.17}_{-4.10}$ $3.79^{+0.37}_{-2.08}$ $\mathbf{2.61^{+0.47}_{-0.42}}$
$m_{3}\sin i$ ($M_{J}$) $0.168^{+0.07}_{-0.07}$ $0.18^{+0.02}_{-0.02}$ $0.33^{+0.10}_{-0.10}$ $0.16^{+0.02}_{-0.02}$ $1.34^{+0.11}_{-0.10}$ $\mathbf{0.15^{+0.01}_{-0.02}}$
$AU3$ $1.995^{+0.25}_{-0.25}$ $1.97^{+0.03}_{-0.03}$ $2.436^{+1.52}_{-1.52}$ $2.00^{+0.05}_{-0.05}$ $0.67^{+1.78}_{-0.67}$ $\mathbf{2.54^{+0.02}_{-0.01}}$
$\gamma_{UCLES}$ (m/s) - - $18.06^{+2.71}_{-1.83}$ - $21.28^{+0.06}_{-1.53}$ $\mathbf{1.78^{+0.37}_{-0.94}}$
$\gamma_{TERRA1}$ (m/s) $1.18^{+2.07}_{-0.19}$ $1.87^{+0.69}_{-0.16}$ $31.28^{+1.79}_{-0.21}$ - - $\mathbf{12.10^{+0.41}_{-0.76}}$
$\gamma_{TERRA2}$ (m/s) $18.20^{+28.31}_{-15.94}$ $14.78^{+2.93}_{-2.50} -$ $3.91^{+0.20}_{-0.91}$ - $\mathbf{-8.373^{+0.74}_{-1.99}}$
$\gamma_{PFS}$ (m/s) - $15.86^{+3.25}_{-2.49} -$ $3.18^{+0.30}_{-0.88}$ - $\mathbf{-8.65^{+0.70}_{-2.41}}$
[…]
$\sigma_{UCLES}$ (m/s) - - $94^{+0.99}_{-0.70} - 4.4$$41^{+0.92}_{-0.63}$ $\mathbf{5.65^{+0.81}_{-0.67}}$
$\sigma_{TERRA1}$ (m/s) $2.03^{+0.45}_{-0.38}$ $1.93^{+0.50}_{-0.37}$ $2.49^{+0.03}_{-0.36}$ - - $\mathbf{1.61^{+0.18}_{-0.13}}$
$\sigma_{TERRA2}$ (m/s) $33^{+0.09}_{-0.05} 1.1$$13^{+0.24}_{-0.14} -$ $1.25^{+0.14}_{-0.13}$ - $\mathbf{1.38^{+0.17}_{-0.15}}$
$\sigma_{PFS}$ (m/s) - $3.15^{+0.58}_{-0.45}$ - $3.66^{+0.10}_{-0.31}$ - $\mathbf{3.12^{+0.29}_{-0.12}}$

### v3

# Destruction of ”peas in a pod?” A candidate multi-planet system around the nearby, bright star, HD208487

@@ Abstract
Aims. We reinvestigate the HD208487 star system to test the reality of the proposed HD208487c world. We also search for additional companions using applied Bayesian statistics and 15+ years of new radial velocity (RV) data from the HARPS and the PFS instruments that were taken post-discovery of HD208487b, as part of our continued study of bright Sun-like stars within 50 pc of the Sun.
Methods. RV data were analyzed with Generalized Lomb-Scargle Periodograms, followed by Bayesian analysis techniques using the EMPEROR code. We also scrutinized various stellar activity indices to search for any corresponding peaks in the power spectra, correlations with the RV measurements, or significant signals from a Bayesian analysis methodology. Finally, photometric data were also checked to test for any transits or possible activity manifestations that could lead to possible false RV signals or excess noise.
@@
Conclusions. We uncovered a candidate three-planet system that would consist of an inner gas giant, a central Saturn, and an outer super-Neptune/sub-Saturn. Extensive analysis of both photometric and spectroscopic data as activity proxies strongly supports the planetary system hypothesis, yet more long-term RV data would help add more statistical weight to the reality of candidate planets c and d. Assuming our model best represents reality, a dynamical analysis suggests that gravitational scattering of an initially ordered, equally spaced system in a long resonant chain of six Neptunes can explain the current architecture of HD208487- a moderately-eccentric, inner massive planet with an outer nearly resonant ($P_{d}/P_{c}=1.495$) gas giant and super-Neptune. More RVs may also shed light on the reality of a fourth Doppler signal uncovered in the data that sits close to the 2:1 period ratio with signal of HD208487c.
@@
Exoplanets – HD208487 – Multiplanetary Systems – Radial Velocity Method – Stellar Activity Modeling
@@ 1 Introduction
In this paper, we will discuss how our new analysis has lead to the statistical confirmation and discovery of two new planets in this system, HD208487c and HD208487d. The rest of the paper will follow the ensuing structure: Sect. 2 will discuss the RV observations in detail, including a breakdown of the measurements by telescope. In Sect. 3 we present updated stellar parameters for HD208487 using the SPECIES and ARIADNE codes. We begin the analysis of the RV dataset by observing the GLS periodograms in Sect. 4, while Sect. 5 details the detection of the three planets with the more rigorous Bayesian analysis. Next, we highlight the search for correlations between stellar activity indices and the RVs in Sect. 6. Sect. 7 describes the analysis of ASAS photometric data and the fitting of stellar cycles. Finally, we discuss the results and their implications in Sect. 8, followed by the conclusions in Sect. 9.
@@ 2.2.2 TERRA2
2.3 6.5m Carnegie Planet Finder Spectrograph (PFS)
@@ 4.2 Split RV data set
As a first check to test whether stellar activity is adding noise to the RV data, or indeed causing the signals we are seeing, we repeated the periodogram analysis with split data sets, where the split is made after calculating the median S-index value in each data set (excluding UCLES data, which do not have S-index values recorded). We then split the RVs in half: one data set with all RVs corresponding to S-index values less than the median and one data set with all RVs corresponding to S-index values greater than or equal to the median. We found that the low S-index data find all three signals, while the high S-index data only find the inner gas giant. Thus, the low S-index data provide a significantly better fit to the data when compared to the high S-index data, which would be expected if the RV signals are all independent of the stellar activity.
@@ 5 Bayesian analysis
After examining the results from the periodogram analyses, we found significant evidence for two additional planetary signals in the data, along with the previously confirmed planet, HD208487b. In order to corroborate the initial results, we followed similar procedures to those employed in Tuomi et al. (2014); Jenkins & Tuomi (2014); D´ıaz et al. (2020) to model the RVs. The Markov Chain Monte Carlo (MCMC) code, Exoplanet Mcmc Parallel tEmpering Radial velOcity fitteR [^9] https://github.com/ReddTea/astroemperor/tree/master (Peña Rojas & Jenkins, 2025), ”searches for signals in radial velocity timeseries, employing Adaptive Parallel Tempering MCMC, convergence tests and Bayesian statistics, along with various noise models.” Further, the Python code provides Bayesian posterior distribution models and the parameterization for the planetary signals to describe the best-fit orbital solution to the star system. An early version of EMPEROR has already been used in a number of works to detect small signals in RV data (Wittenmyer et al., 2017; Jenkins et al., 2019a, b; Barnes et al., 2020; Jenkins et al., 2020), where the updated version employs a custom-written Adaptive Parallel Tempering MCMC engine, similar to the EMCEE package (Foreman-Mackey et al., 2013), as used in (Gregory, 2005, 2007).
@@ 5.2 Model selection
Additionally, we ran each of these models with a fixed eccentricity of zero, a tightly constrained eccentricity described by a normal distribution centered at zero with a standard deviation of 0.1, and a constrained eccentricity described by a normal distribution centered at zero with a standard deviation of 0.3 (as explained in Jenkins et al. (2025)), which enables EMPEROR to explore a wide variety of solutions based on the potential planetary eccentricities. In general, we found that fixing the eccentricity prior to zero would cause additional signals to appear that were harmonics of the real signals, primarily a consequence of the fact that the largest signal in the data corresponds to a massive planet on a non-circular orbit, and hence is poorly explained by a strict cosine-like function. All circular models that caused the 1st harmonic of this primary signal (HD208487b) to appear as a secondary signal in the data performed statistically much worse, enabling us to securely disregard these harmonic models. On the other hand, most of the models with the constrained eccentricity prior or the tightly constrained eccentricity prior yield very similar results. The main difference between the constrained and tightly constrained models are the predicted eccentricities of the outer two planets. The periods and semi-amplitudes are similar, but the constrained models tend to terminate with higher eccentricities, which results in both worse final statistics and long-term unstable planetary dynamics. In the theoretical limit of infinite walkers and steps, we would expect all of the models to converge to the same parameter configuration, even though we extended our analyses as far as reasonably possible to 288 million iterations.
@@ 5.2.2 Comparison of earlier data and later data
$P$ (days) $129.33^{+0.04}_{-0.04}$ $129.32^{+0.03}_{-0.01}$ $129.39^{+0.15}_{-0.22}$ $129.34^{+0.01}_{-0.03}$ $129.27^{+0.23}_{-0.14}$ $129.36^{+0.02}_{-0.02}$
$K$ (m/s) $18.38^{+0.08}_{-0.23}$ $18.76^{+0.27}_{-0.35}$ $17.35^{+0.95}_{-1.22}$ $18.53^{+0.10}_{-0.17}$ $16.70^{+0.95}_{-0.02}$ $18.28^{+0.29}_{-0.40}$
$e$ $0.36^{+0.01}_{-0.02}$ $0.35^{+0.01}_{-0.01}$ $0.18^{+0.04}_{-0.06}$ $0.35^{+0.00}_{-0.01}$ $0.13^{+0.03}_{-0.02}$ $0.37^{+0.01}_{-0.01}$
$\omega$ (rad) $1.21^{+0.03}_{-0.01}$ $1.21^{+0.01}_{-0.03}$ $41^{+0.55}_{-0.35} 1.1$$16^{+0.06}_{-0.04}$ $2.58^{+0.48}_{-0.65}$ $1.10^{+0.01}_{-0.01}$
$M_{0}$ (rad) $3.43^{+0.08}_{-0.10}$ $3.40^{+0.07}_{-0.01}$ $1.83^{+0.38}_{-0.68}$ $3.23^{+0.04}_{-0.05}$ $0.58^{+0.78}_{-0.49}$ $1.95^{+0.05}_{-0.04}$
$m\sin i$ ($M_{J}$) $0.47^{+0.01}_{-0.01}$ $0.48^{+0.01}_{-0.01}$ $0.47^{+0.03}_{-0.03}$ $0.47^{+0.01}_{-0.01}$ $0.45^{+0.03}_{-0.00}$ $0.46^{+0.01}_{-0.00}$
$AU1$ $0.52^{+0.00}_{-0.00}$ $0.52^{+0.00}_{-0.00}$ $0.53^{+0.00}_{-0.00}$ $0.53^{+0.00}_{-0.00}$ $0.52^{+0.00}_{-0.00}$ $0.53^{+0.00}_{-0.00}$
$P_{2}$ (days) $490.82^{+5.62}_{-0.31}$ $488.10^{+0.89}_{-1.54}$ $934.56^{+17.97}_{-42.29}$ $485.60^{+4.71}_{-6.68}$ $923.92^{+26.38}_{-19.30}$ $923.06^{+2.02}_{-2.76}$
$K_{2}$ (m/s) $3.73^{+0.27}_{-0.01}$ $4.35^{+0.33}_{-0.46}$ $9.76^{+1.26}_{-1.98}$ $4.03^{+0.17}_{-0.09}$ $8.69^{+1.13}_{-0.21}$ $6.18^{+0.09}_{-0.07}$
$e_{2}$ $0.17^{+0.02}_{-0.06}$ $0.30^{+0.06}_{-0.10}$ $0.17^{+0.06}_{-0.12}$ $22^{+0.01}_{-0.04} 0.0$$08^{+0.03}_{-0.03}$ $0.19^{+0.01}_{-0.02}$
$\omega_{2}$ (rad) $4.55^{+0.01}_{-0.40}$ $4.72^{+0.33}_{-0.48}$ $3.89^{+1.20}_{-0.04}$ $4.00^{+0.14}_{-0.17}$ $2.94^{+1.46}_{-1.10}$ $5.96^{+0.30}_{-5.96}$
$M_{02}$ (rad) $0.74^{+0.90}_{-0.98}$ $0.41^{+0.19}_{-0.14}$ $3.61^{+1.34}_{-1.70}$ $2.39^{+0.00}_{-0.32}$ $3.96^{+0.00}_{-1.07}$ $1.29^{+0.52}_{-0.29}$
$m_{2}\sin i$ ($M_{J}$) $0.16^{+0.01}_{-0.01}$ $0.18^{+0.01}_{-0.01}$ $0.51^{+0.07}_{-0.07}$ $0.17^{+0.01}_{-0.01}$ $0.46^{+0.06}_{-0.01}$ $0.32^{+0.01}_{-0.01}$
$AU2$ $1.28^{+0.01}_{-0.01}$ $1.27^{+0.01}_{-0.01}$ $1.96^{+0.03}_{-0.03}$ $1.27^{+0.01}_{-0.01}$ $1.95^{+0.04}_{-0.03}$ $1.94^{+0.01}_{-0.01}$
$P_{3}$ (days) $959.17^{+419.38}_{-194.63}$ $941.38^{+28.27}_{-18.51}$ $1294.65^{+1191.79}_{-38.62}$ $979.74^{+33.79}_{-36.32}$ $186.57^{+1125.53}_{-186.57}$ $1380.13^{+19.20}_{-8.25}$
$K_{3}$ (m/s) $3.16^{+1.01}_{-0.94}$ $3.54^{+0.36}_{-0.50}$ $5.61^{+1.63}_{-3.77}$ $3.09^{+0.42}_{-0.37}$ $4.35^{+0.07}_{-1.57}$ $2.46^{+0.30}_{-0.09}$
$e_{3}$ $0.03^{+0.06}_{-0.05}$ $22^{+0.08}_{-0.17} 0.0$$6^{+0.03}_{-0.02} 0.0$$7^{+0.04}_{-0.03} 0.0$$^{+0.08}_{-0.01} 0.110$$1^{+0.03}_{-0.06}$
$\omega_{3}$ (rad) $1.08^{+2.25}_{-1.69}$ $0.57^{+2.24}_{-1.65}$ $2.80^{+1.49}_{-0.65}$ $3.34^{+0.12}_{-1.12}$ $4.91^{+1.86}_{-3.09}$ $0.29^{+0.28}_{-0.09}$
$M_{03}$ (rad) $3.30^{+0.80}_{-1.37}$ $3.27^{+0.37}_{-0.82}$ $5.47^{+1.85}_{-3.32}$ $5.67^{+2.17}_{-4.10}$ $3.79^{+0.37}_{-2.08}$ $2.61^{+0.47}_{-0.42}$
$m_{3}\sin i$ ($M_{J}$) $0.168^{+0.07}_{-0.07}$ $0.18^{+0.02}_{-0.02}$ $0.33^{+0.10}_{-0.10}$ $0.16^{+0.02}_{-0.02}$ $1.34^{+0.11}_{-0.10}$ $0.15^{+0.01}_{-0.02}$
$AU3$ $1.995^{+0.25}_{-0.25}$ $1.97^{+0.03}_{-0.03}$ $2.436^{+1.52}_{-1.52}$ $2.00^{+0.05}_{-0.05}$ $0.67^{+1.78}_{-0.67}$ $2.54^{+0.02}_{-0.01}$
$\gamma_{UCLES}$ (m/s) - - $18.06^{+2.71}_{-1.83}$ - $21.28^{+0.06}_{-1.53}$ $1.78^{+0.37}_{-0.94}$
$\gamma_{TERRA1}$ (m/s) $1.18^{+2.07}_{-0.19}$ $1.87^{+0.69}_{-0.16}$ $31.28^{+1.79}_{-0.21}$ - - $12.10^{+0.41}_{-0.76}$
$\gamma_{TERRA2}$ (m/s) $18.20^{+28.31}_{-15.94}$ $14.78^{+2.93}_{-2.50} -$ $3.91^{+0.20}_{-0.91}$ - $8.373^{+0.74}_{-1.99}$
$\gamma_{PFS}$ (m/s) - $15.86^{+3.25}_{-2.49} -$ $3.18^{+0.30}_{-0.88}$ - $8.65^{+0.70}_{-2.41}$
$\dot{\gamma}$ (m/s yr) $^{+0.00}_{-0.01} 0.000.00$$^{+0.00}_{-0.00}$ $^{+0.00}_{-0.00} 0.000.00$$^{+0.00}_{-0.00}$ $^{+0.00}_{-0.00} 0.000.00$$^{+0.00}_{-0.00}$
$\sigma_{UCLES}$ (m/s) - - $94^{+0.99}_{-0.70} - 4.4$$41^{+0.92}_{-0.63}$ $5.65^{+0.81}_{-0.67}$
$\sigma_{TERRA1}$ (m/s) $2.03^{+0.45}_{-0.38}$ $1.93^{+0.50}_{-0.37}$ $2.49^{+0.03}_{-0.36}$ - - $1.61^{+0.18}_{-0.13}$
$\sigma_{TERRA2}$ (m/s) $33^{+0.09}_{-0.05} 1.1$$13^{+0.24}_{-0.14} -$ $1.25^{+0.14}_{-0.13}$ - $1.38^{+0.17}_{-0.15}$
$\sigma_{PFS}$ (m/s) - $3.15^{+0.58}_{-0.45}$ - $3.66^{+0.10}_{-0.31}$ - $3.12^{+0.29}_{-0.12}$